OpenStudy (anonymous):

mc094-1.jpg and mc094-2.jpgare complementary angles. mc094-3.jpg = x + 5 and mc094-4.jpg = x – 9. The measure of mc094-5.jpg is _____, and the measure of mc094-6.jpg is _____. 47°, 43° 52°, 48° 52°, 38° 47°, 53° 14. mc090-1.jpg How are mc090-2.jpg and mc090-3.jpg related? supplementary angles vertical angles adjacent angles complementary angles 15. mc022-1.jpg When this net is folded into a cube, which two letters appear on opposite sides? B and E C and D D and A A and E 16. mc036-1.jpg This net folds into the cube shown beside it. On the cube,

OpenStudy (texaschic101):

I cannot see the pictures....but I think I can figure one of them out.. complimentary angles add up to 90 degrees... x + 5 + x - 9 = 90 2x - 4 = 90 2x = 94 x = 47 x + y = 90 47 + y = 90 y = 90 - 47 y = 43 so your angles are : 47 and 43 ============== need the pictures for the rest of them...